1 Rutherford Rutherford discovered the nucleus of the atom by firing a particles at gold foil. An a particle has a charge of q+2e and a mass of m6.64 10-27 kg. A gold nucleus has charge of +79e. You may ignore the motion of the gold nucleus in this problem +2e +79e Suppose an a particles is traveling directly toward a gold nucleus. If the speed of the particle is v-19-107 m/s when it is 1 m from the gold nucleus, how close to the gold nucleus will the a particle come before it stops and reverses direction?
Answer #1

By conservation of energy

1/2mv^2+kq1q2/r=kq1q2/r'

k=9*10^9Nm^2/C^2

Substituting values

m=6.64*10^-27kg

q1=2e

q2=79e

v=1.9*10^7m/s

r=1m

we get r'=303.866*10^-14m